OpenStudy (anonymous):

Find the product of the complex number and it's conjugate; 7-6t

jimthompson5910 (jim_thompson5910):

hint: (a+bi)*(a-bi) = a^2 + b^2

OpenStudy (jdoe0001):

7-6t doesn't look very complex per se

OpenStudy (anonymous):

I'm sorry I really don't understand

jimthompson5910 (jim_thompson5910):

the number is 7-6i right?

OpenStudy (anonymous):

Yes, but how do I get a product from that

jimthompson5910 (jim_thompson5910):

the conjugate is 7+6i

jimthompson5910 (jim_thompson5910):

multiply the two numbers using the rule I posted above

jimthompson5910 (jim_thompson5910):

or you can expand it out the long way

OpenStudy (anonymous):

How do I get a product withput an = sign?

jimthompson5910 (jim_thompson5910):

in this case, a = 7 and b = 6

jimthompson5910 (jim_thompson5910):

so... (a+bi)*(a-bi) = a^2 + b^2 (7+6i)*(7-6i) = 7^2 + 6^2 (7+6i)*(7-6i) = ???

OpenStudy (anonymous):

THANK YOU SO MUCH

jimthompson5910 (jim_thompson5910):

yw